OpenStudy (anonymous):

Evaluate the expression 4a^2 - 2b / 4 for a=2 and b=4. 2 8 14 16

OpenStudy (shamil98):

input your values of a and b and evaluate it then..

OpenStudy (anonymous):

substitute the values of a & b in the equation , and then all you have will be numeric ordinary expression

OpenStudy (anonymous):

The answer is C ?

OpenStudy (anonymous):

like 4 * (2)^2 = 4 * 4 =16 then 2b/4 = (2)(4)/4 =8 /4 = 2 then you'll have 16 - 2 = 14

OpenStudy (shamil98):

Do not guess. Do the work, and show it here.

OpenStudy (shamil98):

4a^2 - 2b / 4 4*4 - 2(4) / 4 8 /4 = 2

OpenStudy (anonymous):

@shamil98 I didn't guess. The answer is C!

OpenStudy (anonymous):

yeah ^^

OpenStudy (shamil98):

is the question \[\huge \frac{ 4a^2 - 2b }{ 4 }\] or \[4a^2 - \frac{ 2b }{ 4 }\]

OpenStudy (shamil98):

if it is the second one then yes it is 14. if it is the first then it is 2.

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

14

OpenStudy (shamil98):

In that case the -2b is divided by 4 only so it would be the second equation i posted. so yes 14.

OpenStudy (anonymous):

Thanks @shamil98